# Benchmark on Drug Target Interaction Modeling from a Structure Perspective

### v1

@@ 4 A Macroscopical Benchmark on Encoder and Featurization Strategies
Encoder Exploration for Drugs and Proteins. To investigate the influence of different encoding strategies for extracting the structural information of drugs, we employ GCN [17] and vanilla Transformer [56] as the encoders for drugs. Meanwhile, integer encoding with CNN, n-gram encoding with CNN, and the vanilla Transformer are considered to capture protein’s representations, which are frequently adopted. To leverage the advantages of the pretrained protein information, we include a language model, i.e., Evolutionary Scale Modeling (ESM2) [49]. The results of various combinations of drug and protein encoders are listed in Table 1.
@@
Obs. 1. GNN and Transformer-based drug encoders exhibit unequal performance depending on DTI tasks. When the encoder for the protein sequence is fixed, drug features extracted by the Transformer generally perform better than those by GNNs in regression tasks, but the opposite is true in classification tasks. Notably, in the classification tasks on Human dataset, the combination of GNN and Transformer used, respectively, for drugs and proteins yields excellent performance but falls short in the regression task. This disparity may be due to the smaller size of the Human dataset compared to the Davis dataset, which allows for faster convergence in classification tasks than in regression tasks under a fixed epoch.
@@
Obs. 2. Transformer models are better in extracting features from protein. Although we only consider the simplest pretrained protein language model of ESM2, it still significantly outperforms other encoders. This improvement can likely be attributed to the robust and generalizable representations learned from extensive data by the pretrained model. In the classification task, transformers achieve the best performance, underscoring their effectiveness in extracting protein sequence features.
@@
| Drug Encoder | Protein Encoder | Regression | Classification | |||||||||||
| MSE | MAE | R2 | PCC | CI | Spearman | ROC-AUC | PR-AUC | Log-AUC | Acc. | Precision | Recall | F1 | ||
| GCN | Int encoding with CNN | 0.3181 | 0.3314 | 0.5263 | 0.7315 | 0.8311 | 0.5906 | 0.9122 | 0.8761 | 0.4452 | 0.9127 | 0.9033 | 0.9301 | 0.9165 |
| GCN | 3-gram pretrained with CNN | 0.3361 | 0.3613 | 0.4994 | 0.7133 | 0.8332 | 0.5956 | 0.9016 | 0.8634 | 0.4045 | 0.9021 | 0.8944 | 0.9183 | 0.9061 |
| GCN | ESM2 | 0.2611 | 0.3112 | 0.6111 | 0.7843 | 0.8665 | 0.6480 | 0.9331 | 0.9046 | 0.5034 | 0.9334 | 0.9275 | 0.9446 | 0.9371 |
| GCN | Trans | 0.2911 | 0.3326 | 0.5664 | 0.7600 | 0.8571 | 0.6329 | 0.9435 | 0.9192 | 0.5282 | 0.9438 | 0.9395 | 0.9524 | 0.9458 |
| Trans | Int Encoding with CNN | 0.2553 | 0.2853 | 0.6197 | 0.7890 | 0.8620 | 0.6397 | 0.9096 | 0.8783 | 0.4623 | 0.9096 | 0.9142 | 0.9099 | 0.9121 |
| Trans | 3-gram pretrained with CNN | 0.2618 | 0.3032 | 0.6101 | 0.7820 | 0.8581 | 0.6338 | 0.9011 | 0.8674 | 0.4327 | 0.9012 | 0.9054 | 0.9030 | 0.9040 |
| Trans | ESM2 | 0.2609 | 0.3052 | 0.6115 | 0.7869 | 0.8617 | 0.6391 | 0.9203 | 0.8924 | 0.4902 | 0.9203 | 0.9257 | 0.9192 | 0.9224 |
| Trans | Trans | 0.2828 | 0.3222 | 0.5788 | 0.7721 | 0.8554 | 0.6297 | 0.9306 | 0.9073 | 0.5158 | 0.9304 | 0.9391 | 0.9250 | 0.9319 |
@@
Featurization Exploration. Despite the efficacy of GNNs in learning drug structures, the featurization of nodes plays a critical role in capturing both the intrinsic properties of atoms and their contextual relevance. We conduct a detailed analysis of various methods (summarized in Section F of Appendix) for constructing graph features within the DTI context. The node feature is constructed via various characteristics, such as chemical and physical properties. We categorize each feature into five main classes, e.g., atomic properties (AP), hydrogen information (HI), electron properties (EP), stereochemistry (Ste) and structural information (Str). To better determine which types of features are more effective in capturing the structural information, we conduct an ablation study on the different featurization strategies. Here we choose GprahDTA [18] and GraphCPI [35] as our backbone models. The results of feature combinations are reported in Fig. 2.
Obs. 4. More complex featurization does not necessarily bring positive effect. Despite employing different protein representations (different colors in Fig. 2), GraphCPI and GraphDTA perform stably on the benchmark dataset Davis. Moreover, the basic feature configuration has the lowest MSE compared to other featurization strategies, suggesting that additional features may complicate the model’s ability to discern critical information for regression task. Additionally, the increased CI with more complex features, such as stereochemistry and structure, suggests that while they might not improve the prediction accuracy, they do contribute positively to the model’s ranking capability.
Obs. 5. The role of atomic and electron properties in modeling drug features may not be inherently detrimental, but their contribution appears to be context-dependent. In Fig. 2, it suggests that these properties, when not combined with other informative features, don’t significantly enhance two models’ performance. In the classification task, as reflected by the F1 scores and ROC-AUC, it is evident that stereochemistry and structural information substantially improve the model’s accuracy. Thus, while atomic and electron properties are fundamental, their full potential is unlocked when integrated with stereochemical and structural information, underscoring the importance of a multifaceted approach in node featurization.
@@ 5 A Microscopic Benchmark on DTI Models
Benchmark over Effectiveness. As shown in Table. 2 and Table. 3, we conduct experiments on 31 different models across two tasks and three datasets, respectively (see more comprehensive model comparisons in Section G of Appendix). All results are averaged by five-fold cross-validation.
Obs. 6. Molecular graphs are better than fingerprints to capture the graph features of drug. In reference to Table. 7, it is evident that GNN-based approaches utilizing the molecular graph generally yield superior performance compared with fingerprints (CPI [31], BACPI [57], GANDTI [58]). This reinforces the idea that the rich structural and atom property information inherent to molecular graphs is pivotal for representation extraction, leading to enhanced model performance.
Obs. 7. Graph structure is a crucial part in extracting drug’s features. Different GNNs have the distinct performances in both tasks when the protein representation is fixed. Specifically, GIN, with its unique ability to distinguish non-isomorphic graphs, consistently outperforms other models across different protein encoders in regression tasks. Although transformer-based methods such as MRBDTA are proficient in handling sequential information from SMILES and proteins, the depth of information they capture appears to be marginally less comprehensive than that provided by molecular graph-based approaches. This is substantiated by the superior performance of GNN-based methods, including MGraphDTA, ColdDTA, and SubMDTA, which suggest that GNN captures intricate structural details more effectively.
| Category | Models | DAVIS | KIBA | BindingDB_Kd | Avg. Reduce | ||||||
@@
| GNN | GraphDTA-GIN [18] | 0.2309 | 0.6562 | 0.8711 | 0.0004 | 0.5544 | 0.7996 | 0.5033 | 0.7259 | 0.8576 | 8.5625 |
| GraphCPI-GIN [35] | 0.2413 | 0.6407 | 0.8671 | 0.0005 | 0.4963 | 0.7808 | 0.5069 | 0.7239 | 0.8572 | 10.2845 | |
| MGraphDTA [59] | 0.2179 | 0.6755 | 0.8820 | 0.0003 | 0.7208 | 0.8649 | 0.4887 | 0.7338 | 0.8649 | 4.9795 | |
| SAGDTA [60] | 0.2656 | 0.6045 | 0.8675 | 0.0039 | -3.3777 | 0.8096 | 0.6590 | 0.6410 | 0.8366 | 27.6575 | |
| EmbedDTI [61] | 0.2561 | 0.6186 | 0.8624 | 0.0007 | 0.2364 | 0.6374 | 0.5095 | 0.7225 | 0.8559 | 12.3450 | |
| DeepGLSTM [29] | 0.2915 | 0.5659 | 0.8476 | 0.0003 | 0.6919 | 0.8480 | 0.5385 | 0.7067 | 0.8529 | 19.1015 | |
| CPI [31] | 0.3503 | 0.4784 | 0.8319 | 0.0003 | 0.6274 | 0.8190 | 0.6962 | 0.6208 | 0.8261 | 35.8330 | |
| BACPI [57] | 0.4036 | 0.3990 | 0.7982 | 0.0006 | 0.5320 | 0.8175 | 0.6468 | 0.6477 | 0.8297 | 36.0894 | |
| DeepNC-HGC [30] | 0.2782 | 0.5857 | 0.8551 | 0.0005 | 0.4983 | 0.7823 | 0.5611 | 0.6944 | 0.8464 | 20.0167 | |
| DeepNC-GEN [30] | 0.2543 | 0.6213 | 0.8634 | 0.0004 | 0.5355 | 0.7981 | 0.5561 | 0.6971 | 0.8496 | 17.1559 | |
| DrugBAN [62] | 0.2391 | 0.6440 | 0.8757 | 0.0004 | 0.5611 | 0.8388 | 0.4485 | 0.7557 | 0.8693 | 2.3692 | |
| GANDTI [58] | 0.3082 | 0.5410 | 0.8414 | 0.0003 | 0.6569 | 0.8342 | 0.6714 | 0.6343 | 0.8322 | 31.4522 | |
| PGraphDTA-CNN [24] | 0.3273 | 0.5126 | 0.8701 | 0.0005 | 0.4833 | 0.7473 | 0.5334 | 0.7095 | 0.8590 | 22.0042 | |
| BridgeDPI [63] | 0.3623 | 0.6477 | 0.8991 | 0.0004 | 0.5686 | 0.7849 | 0.4482 | 0.7559 | 0.8698 | 17.1661 | |
| ColdDTA [64] | 0.2346 | 0.6507 | 0.8693 | 0.0004 | 0.5948 | 0.8018 | 0.4697 | 0.7442 | 0.8644 | 4.6828 | |
| SubMDTA [36] | 0.2326 | 0.6537 | 0.8691 | 0.0003 | 0.6855 | 0.8485 | 0.4566 | 0.7513 | 0.8670 | 2.5816 | |
| IMAEN [65] | 0.2412 | 0.6409 | 0.8721 | 0.0004 | 0.5800 | 0.8061 | 0.4720 | 0.7429 | 0.8553 | 5.8716 | |
| Transformer | CSDTI [32] | 0.3029 | 0.5490 | 0.8395 | 0.0007 | 0.2448 | 0.6475 | 0.6408 | 0.6510 | 0.8369 | 28.8755 |
| TDGraphDTA [25] | 0.2217 | 0.6698 | 0.6685 | 0.0008 | 0.0533 | 0.3429 | 0.4750 | 0.7413 | 0.7887 | 3.6989 | |
| AMMVF [38] | 0.3325 | 0.5048 | 0.8307 | 0.0006 | 0.4696 | 0.7711 | 0.6597 | 0.6407 | 0.8336 | 32.3429 | |
| IIFDTI [66] | 0.2741 | 0.5918 | 0.8500 | 0.0005 | 0.2584 | 0.7952 | 0.5097 | 0.7170 | 0.8585 | 14.3568 | |
| ICAN [42] | 0.3481 | 0.4816 | 0.8211 | 0.0008 | 0.1159 | 0.8256 | 0.6582 | 0.6415 | 0.8277 | 33.3035 | |
| MolTrans [20] | 0.2588 | 0.6146 | 0.8601 | 0.0003 | 0.6378 | 0.8453 | 0.5138 | 0.7201 | 0.8570 | 13.0935 | |
| TransformerCPI [21] | 0.2869 | 0.5728 | 0.8326 | 0.0008 | 0.0728 | 0.8357 | 0.5704 | 0.6894 | 0.8426 | 21.7224 | |
| MRBDTA [55] | 0.2350 | 0.6499 | 0.8775 | 0.0006 | 0.3210 | 0.7239 | 0.4977 | 0.7289 | 0.8629 | 8.4004 | |
| FOTFCPI [41] | 0.2803 | 0.5825 | 0.8546 | 0.0004 | 0.5342 | 0.7947 | 0.5743 | 0.6872 | 0.8444 | 21.4386 | |
| Our combos | 0.2063 | 0.6927 | 0.8901 | 0.0003 | 0.7168 | 0.8677 | 0.4651 | 0.7467 | 0.8683 | 0 | |
| Category | Models | Human | C.elegans | Drugbank | Avg. Improve | ||||||
@@
| GNN | GraphDTA-GATGCN[18] | 0.9296 | 0.9297 | 0.9319 | 0.9487 | 0.9484 | 0.9485 | 0.7712 | 0.7711 | 0.7764 | 2.5140 |
| GraphCPI-GATGCN[35] | 0.9097 | 0.9099 | 0.9129 | 0.9372 | 0.9370 | 0.9369 | 0.7561 | 0.7561 | 0.7593 | 4.3335 | |
| MGraphDTA[59] | 0.9408 | 0.9410 | 0.9429 | 0.9631 | 0.9630 | 0.9628 | 0.8146 | 0.8146 | 0.8157 | -0.1030 | |
| SAGDTA[60] | 0.9021 | 0.9018 | 0.9032 | 0.9380 | 0.9378 | 0.9375 | 0.7655 | 0.7655 | 0.7662 | 4.2494 | |
| EmbedDTI[61] | - | - | - | - | - | - | 0.7625 | 0.7625 | 0.7674 | 5.3770 | |
| DeepGLSTM[29] | 0.9180 | 0.9183 | 0.9212 | 0.9414 | 0.9412 | 0.9409 | 0.7640 | 0.7640 | 0.7674 | 3.5182 | |
| CPI[31] | 0.9110 | 0.9111 | 0.9138 | 0.9312 | 0.9310 | 0.9309 | 0.7467 | 0.7466 | 0.7515 | 4.9098 | |
| BACPI[57] | 0.9249 | 0.9251 | 0.9276 | 0.9556 | 0.9555 | 0.9551 | 0.7748 | 0.7748 | 0.7776 | 2.2746 | |
| DeepNC-HGC[30] | 0.8796 | 0.8798 | 0.8834 | 0.9418 | 0.9417 | 0.9413 | 0.7673 | 0.7673 | 0.7718 | 4.9057 | |
| DeepNC-GEN[30] | 0.9178 | 0.9180 | 0.9205 | 0.9501 | 0.9499 | 0.9498 | 0.7464 | 0.7463 | 0.7571 | 3.8865 | |
| DrugBAN[62] | 0.9302 | 0.9302 | 0.9320 | 0.9596 | 0.9594 | 0.9593 | 0.8188 | 0.8188 | 0.8179 | 0.2732 | |
| GANDTI [58] | 0.9333 | 0.9333 | 0.9351 | 0.9404 | 0.9403 | 0.9400 | 0.7427 | 0.7425 | 0.7550 | 3.8110 | |
| PGraphDTA-CNN [24] | - | - | - | - | - | - | 0.7910 | 0.7910 | 0.7948 | 1.5803 | |
| BridgeDPI [63] | 0.9456 | 0.9456 | 0.9471 | 0.9651 | 0.9651 | 0.9646 | 0.7825 | 0.7825 | 0.7853 | 0.8392 | |
| ColdDTA [64] | 0.9420 | 0.9421 | 0.9436 | 0.9623 | 0.9622 | 0.9619 | 0.8186 | 0.8186 | 0.8203 | -0.2608 | |
| SubMDTA [36] | 0.9326 | 0.9328 | 0.9351 | 0.9570 | 0.9568 | 0.9566 | 0.8045 | 0.8045 | 0.8072 | 0.8055 | |
| IMAEN [65] | 0.9058 | 0.9059 | 0.9089 | 0.9506 | 0.9505 | 0.9500 | 0.7995 | 0.7994 | 0.8014 | 2.2592 | |
| Transformer | CSDTI [32] | 0.8630 | 0.8630 | 0.8663 | 0.8962 | 0.8958 | 0.8966 | 0.7269 | 0.7269 | 0.7306 | 9.2569 |
| TDGraphDTA [25] | 0.9411 | 0.9409 | 0.9419 | 0.9573 | 0.9571 | 0.9569 | 0.8083 | 0.8082 | 0.8116 | 0.3547 | |
| AMMVF [38] | 0.9287 | 0.9290 | 0.9314 | 0.9636 | 0.9635 | 0.9632 | 0.7814 | 0.7814 | 0.7849 | 1.5670 | |
| IIFDTI [66] | 0.9392 | 0.9392 | 0.9409 | 0.9679 | 0.9679 | 0.9675 | 0.8084 | 0.8084 | 0.8125 | 0.0110 | |
| ICAN [42] | 0.9387 | 0.9389 | 0.9410 | 0.9644 | 0.9644 | 0.9637 | 0.7731 | 0.7731 | 0.7745 | 1.4721 | |
| MolTrans [20] | 0.9453 | 0.9455 | 0.9471 | 0.9639 | 0.9638 | 0.9634 | 0.7910 | 0.7909 | 0.7922 | 0.5777 | |
| TransformerCPI [21] | 0.9311 | 0.9312 | 0.9333 | 0.9581 | 0.9580 | 0.9577 | 0.8097 | 0.8097 | 0.8125 | 0.6262 | |
| MRBDTA [55] | 0.9447 | 0.9446 | 0.9458 | 0.9713 | 0.9713 | 0.9709 | 0.8102 | 0.8102 | 0.8134 | -0.3778 | |
| FOTFCPI [41] | 0.9444 | 0.9444 | 0.9459 | 0.9673 | 0.9672 | 0.9669 | 0.7845 | 0.7845 | 0.7889 | 0.7307 | |
| Our combos | 0.9435 | 0.9435 | 0.9451 | 0.9688 | 0.9688 | 0.9684 | 0.8035 | 0.8035 | 0.8061 | 0 | |
@@
Benchmark over Convergence. We select the top two methods from the GNN-based and Transformer-based frameworks, respectively, and evaluate them across six datasets on two tasks. The training losses are depicted in Fig. 4. In order to compare different methods, we only show the epochs before 300. Based on the empirical data, we summarize our primary observations as follows:
Obs. 9. GNN-based methods demonstrate quicker and more stable convergence compared to Transformer-based methods. This phenomenon arises from the fact that GNN-based methods have fewer memory usage and model parameters, leading to a larger batch size usage or faster convergence compared with Transformer-based methods.

### v2

@@ 4 A Macroscopic Benchmark on Encoder and Featurization Strategies
Encoder Exploration for Drugs and Proteins. To investigate the influence of different encoding strategies for extracting the structural information of drugs, we employ GIN (gin) and vanilla Transformer (vaswani2017attention) as the encoders for drugs. Meanwhile, integer encoding with CNN, n-gram encoding with CNN, and the vanilla Transformer are considered to capture protein’s representations, which are frequently adopted. To leverage the advantages of the pretrained protein information, we include a language model, i.e., Evolutionary Scale Modeling (ESM2) (lin2022language). The results of various combinations of drug and protein encoders are shown in Fig. 1.All results are averaged by five-fold cross-validation with an early stop mechanism.
@@
Obs. 1. GNN and Transformer-based drug encoders exhibit unequal performance depending on DTI tasks. When the encoder for the protein sequence is fixed, drug features extracted by the GNN structures GIN generally perform better than those by Transformers in regression tasks, but the opposite is true in classification tasks. This disparity may be due to the smaller size of the Human dataset compared to the Davis dataset, which allows for faster convergence in classification tasks than in regression tasks.
@@
Obs. 2. Transformer models are better but sensitive in extracting features from protein. Although we only consider the simplest pretrained protein language model of ESM2, it still significantly outperforms other encoders in both tasks. This improvement can likely be attributed to the robust and generalizable representations learned from extensive data by the pretrained model. In addition, the Transformer encoder for the protein achieves the best performance on the classification task but shows unstable performance in the regression task. This is likely due to the smaller size and simpler classification dataset compared to the regression dataset, making the training stop for a fixed early stop threshold.
@@
Featurization Exploration. Despite the efficacy of GNNs in learning drug structures, the featurization of nodes plays a critical role in capturing both the intrinsic properties of atoms and their contextual relevance. We conduct a detailed analysis of various methods (summarized in Section G of the Appendix) for constructing graph features within the DTI context. The node feature is constructed via various characteristics, such as chemical and physical properties. We categorize each feature into five main classes, e.g., atomic properties (AP), hydrogen information (HI), electron properties (EP), stereochemistry (Ste), and structural information (Str). To better determine which types of features are more effective in capturing the structural information, we conduct an ablation study on the different featurization strategies. We choose GraphDTA (nguyen2021graphdta) and GraphCPI (quan2019graphcpi) with GIN as our backbone models. The results of feature combinations are reported in Fig. 2.
Obs. 4. More complex featurization does not necessarily bring a positive effect, and its effectiveness is highly task-dependent. As shown in Fig. 2, adding features like atomic properties (AP), hydrogen information (HI), and stereochemistry information (Ste) improves performance in the regression task by reducing the MSE loss, suggesting that these features provide valuable information. However, features like electron properties (EP) and structural information (Str) may introduce noise rather than useful information, especially when combined with other features, leading to inconsistent results. Furthermore, in classification tasks, the trend differs, with additional features sometimes negatively impacting performance (blue line) rather than providing benefits. This highlights the importance of careful feature selection, as indiscriminate inclusion of complex features may lead to increased noise, affecting model generalization and robustness.
Obs. 5. Protein representation plays a crucial role in the effectiveness of different drug featurizations. As shown in Fig. 2, even when using the same drug featurization strategies, the trends vary depending on the protein representation. This suggests that the way proteins are encoded directly influences how drug features interact with the model. For instance, in GraphDTA, features like atomic properties (AP) and hydrogen information (HI) help improve performance, but in GraphCPI, those same features don’t always provide the same benefits. In some cases, adding more drug features introduces noise rather than useful information. This highlights the fact that drug featurization and protein representation are deeply interconnected, and optimizing one without considering the other may not yield the best results. To build more effective drug-protein interaction models, both components should be considered holistically rather than in isolation.
@@ 5 A Microscopic Benchmark on DTI Models
Benchmark over Effectiveness. As shown in Table 1 and Table 2, we conduct experiments on models with their optimal hyperparameters across two tasks and three datasets, respectively (see more comprehensive model comparisons in Appendix F). All results are averaged by five-fold cross-validation with an early stop mechanism.
Obs. 6. Molecular graphs are better than fingerprints for capturing the graph features of a drug. In reference to Table 7, it is evident that GNN-based approaches utilizing the molecular graph generally yield superior performance compared with fingerprints (CPI (tsubaki2019compound), BACPI (li2022bacpi), GANDTI (wang2021gandti)). This reinforces the idea that the rich structural and atomic property information inherent to molecular graphs is pivotal for representation extraction, leading to enhanced model performance.
Obs. 7. Graph structure is a crucial part of extracting a drug’s features. Different GNNs have distinct performances in both tasks when the protein representation is fixed. Specifically, GIN, with its unique ability to distinguish non-isomorphic graphs, consistently outperforms other models across different protein encoders in regression tasks. Although Transformer-based methods such as MRBDTA are proficient in handling sequential information from SMILES and proteins, the depth of information they capture appears to be marginally less comprehensive than that provided by molecular graph-based approaches. This is substantiated by the superior performance of GNN-based methods, including MGraphDTA, ColdDTA, and SubMDTA, which suggests that GNN captures intricate structural details more effectively.
| Category | Models | DAVIS | KIBA | BindingDB_Kd | Avg. Reduction | ||||||
@@
| GNN | GraphDTA-GIN | ||||||||||
| GraphCPI-GIN | |||||||||||
| MGraphDTA | |||||||||||
| SAGDTA | |||||||||||
| EmbedDTI | |||||||||||
| DeepGLSTM | |||||||||||
| CPI | |||||||||||
| BACPI | |||||||||||
| DeepNC-HGC | |||||||||||
| DeepNC-GEN | |||||||||||
| DrugBAN | |||||||||||
| GANDTI | |||||||||||
| BridgeDPI | |||||||||||
| ColdDTA | |||||||||||
| SubMDTA | |||||||||||
| IMAEN | |||||||||||
| Transformer | CSDTI | ||||||||||
| TDGraphDTA | |||||||||||
| AMMVF | |||||||||||
| IIFDTI | |||||||||||
| ICAN | |||||||||||
| MolTrans | |||||||||||
| TransformerCPI | |||||||||||
| MRBDTA | |||||||||||
| FOTFCPI | |||||||||||
| Our combos | |||||||||||
| Category | Models | Human | C.elegans | Drugbank | Avg. Improvement | ||||||
@@
| GNN | GraphDTA-GIN | 3.312% | |||||||||
| GraphCPI-GIN | 4.275% | ||||||||||
| MGraphDTA | 1.322% | ||||||||||
| SAGDTA | 4.600% | ||||||||||
| EmbedDTI | 3.871% | ||||||||||
| DeepGLSTM | 4.356% | ||||||||||
| CPI | 7.708% | ||||||||||
| BACPI | 2.522% | ||||||||||
| DeepNC-HGC | 6.006% | ||||||||||
| DeepNC-GEN | 4.194% | ||||||||||
| DrugBAN | 0.675% | ||||||||||
| GANDTI | 3.671% | ||||||||||
| BridgeDPI | 11.189% | ||||||||||
| ColdDTA | 0.037% | ||||||||||
| SubMDTA | 1.055% | ||||||||||
| IMAEN | 4.560% | ||||||||||
| Transformer | CSDTI | 11.467% | |||||||||
| TDGraphDTA | 0.299% | ||||||||||
| AMMVF | 6.595% | ||||||||||
| IIFDTI | 1.437% | ||||||||||
| ICAN | 1.899% | ||||||||||
| MolTrans | 0.525% | ||||||||||
| TransformerCPI | 0.979% | ||||||||||
| MRBDTA | 0.789% | ||||||||||
| FOTFCPI | 0.112% | ||||||||||
| Our combos | 0.000% | ||||||||||
@@
Benchmark over Convergence. We select the two representative methods from the GNN-based and Transformer-based frameworks, respectively, and evaluate them across six datasets on two tasks. The training losses are depicted in Fig. 4. To ensure a fair comparison of convergence behavior, we use the previous early stopping setting. Based on the empirical results, we summarize our key observations as follows:
Obs. 9. GNN-based methods demonstrate quicker convergence compared to Transformer-based methods. This phenomenon arises from the fact that GNN-based methods have less memory usage and fewer model parameters, leading to larger batch size usage or faster convergence compared with Transformer-based methods.